E18-23 (similar to)

Johnson

Winery in Pleasant​ Valley, New​ York, has two​ departments: Fermenting and Packaging. Direct materials are added at the beginning of the fermenting process​ (grapes) and at the end of the packaging process​ (bottles). Conversion costs are added evenly throughout each process. The company uses the​ weighted-average method. Data from the month of March for the Fermenting Department are as​ follows:

Gallons

Beginning Work-in-Process Inventory

400

gallons

Started in production

7,500

gallons

Completed and transferred out to Packaging in March

4,400

gallons

Ending Work-in-Process Inventory (80% of the way

3,500

gallons

through the fermenting process)

Costs

Beginning Work-in-Process Inventory:

Direct materials

$1,264

Direct labor

160

1.

Compute the Fermenting​ Department's equivalent units of production for direct materials and for conversion costs.

2.

Compute the total costs of the units​ (gallons)

a. completed and transferred out to the Packaging Department.

b. in the Fermenting Department ending​ Work-in-Process Inventory.

Requirement 1. Compute the Fermenting​ Department's equivalent units of production for direct materials and for conversion costs.

Complete the partial production cost report below for the Fermenting​ Department, showing the equivalent units of production for direct materials and for conversion costs.

Johnson Winery

Production Cost Report - Fermenting Department (Partial)

Month Ended March 31

Equivalent Units

Physical

Direct

Conversion

UNITS

Units

Materials

Costs

Units to account for:

Started in production

Total units to account for

Units accounted for:

Total units accounted for

Choose from any list or enter any number in the input fields and then click Check Answer.
Answer #1

Answer

Johnson Winery

Production Cost Report - Fermenting Department (Partial)

Month Ended March 31

Equivalent Units

Physical

Direct

Conversion

UNITS

Units

Materials

Costs

Units to account for:

Beginning WIP Inventory

                           400

Started in production

                       7,500

Total units to account for

                       7,900

Units accounted for:

Units completed & transferred out

                       4,400

               4,400

                  4,400

Unit in ending WIP inventory

                       3,500

               3,500

                  2,800

Total units accounted for

                       7,900

               7,900

                  7,200

--Working

Johnson Winery

Production Cost Report - Fermenting Department (Partial)

Month Ended March 31

Equivalent Units

Physical

Direct

Conversion

UNITS

Units

Materials

Costs

Units to account for:

Beginning WIP Inventory

400

Started in production

7500

Total units to account for

=400+7500

Units accounted for:

Units completed & transferred out

4400

=4400*100%

=4400*100%

Unit in ending WIP inventory

3500

=3500*100%

=3500*80%

Total units accounted for